OpenStudy (anonymous):

last one... xy^2 --------- please help! 6x^-2y^-2

OpenStudy (anonymous):

\[\frac{x^3y^4}{6}\]

OpenStudy (anonymous):

Then move the negative exponents to the top

OpenStudy (anonymous):

whattt???

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

\[\frac{x^{1-(-2)}y^{2-(-2)}}{6}\]

OpenStudy (anonymous):

Which gives you \[\frac{x^3y^4}{6}\]

OpenStudy (anonymous):

Sorry i got confused for a minute lol

OpenStudy (anonymous):

ah i get it(:

OpenStudy (anonymous):

Yeah, i didnt like that one lol :P